(2x-2)+(3x-3)+(9x-9) = 1
 (2x-2)+(3x-3)+(9x-9) = 1
 (2x-2)+(3x-3)+(9x-9) = 1
 (2x-2)+(3x-3)+(9x-9) = 1

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(2x-2)+(3x-3)+(9x-9) = 1

2x – 2 + 3x – 3 + 9x – 9 = 1

14x – 14 = 1

14x = 1 + 14

x = 15/14

Thanks